Rearrange the following equation to find x in terms of a and b, in
simplest form (assume that a '= 4b):
a(a − x) + 4b(x − 4b) = 0.
This is my solution am i on the right track:
a^2-ax+4bx-16b^2=0
a^2-ax-0=4bx-16b^2
a^2-ax=4bx-16b^2
Now do i set each one to equal zero? on the algebra calculator i can see that the answer is x=4b+a but i have no idea how to get this answer. Thanks

Answer by oscargut(2103)

a(a − x) + 4b(x − 4b) = 0.
This is my solution am i on the right track:
a^2-ax+4bx-16b^2=0 (right)
(4b-a)x = 16b^2-a^2

(4b-a)x = (4b-a)(4b+a) , since a is not equal to 4b (4b-a is not 0)
x = 4b+a
If you need more help with similar questions, homework, assignments, you can contact me now at:
mthman@gmail.com
Thanks